OpenStudy (anonymous):

In a classroom, 8 out of 19 students surveyed say they had two or more siblings. If the entire school has 570 students, predict the number of students that have two or more siblings.

OpenStudy (campbell_st):

the number is \[\frac{8}{19} \times 570 =\] then round the answer to the nearest whole number

OpenStudy (anonymous):

Use simple proportion: 8...........19 x...........570 x=(570*8)/19